OpenStudy (anonymous):

supplementary angles are angles whose measures have a sum of 180degree complementary angles have a sum of 90degrees find the measure of an angle whose supplement is 10degrease more than twice its supplement let 90-x equal the degree measure of its supplement write and solve and equation!

OpenStudy (anonymous):

180-x=10+2(90-x) 180-x=10+180-2x x=10 complementary=80 supplementary=170=10+2*80